OpenStudy (kitkat16):

Please help my brain is fried. :( David rowed a boat upstream for three miles and then returned to point he started from. The entire journey took four hours. The speed of the stream is one mile per hour. Find David's speed in still water. David's speed in still water is ? miles per hour?

OpenStudy (kitkat16):

speed = distance ÷ time, upstream speed = speed of the boat – speed of the stream, and downstream speed = speed of the boat + speed of the stream

OpenStudy (kitkat16):

@iGreen u busy?

OpenStudy (kitkat16):

@Directrix any ideas

OpenStudy (kitkat16):

s=6 divided by 4 =3/2=1.5 s=1.5

OpenStudy (kitkat16):

upstream speed =1.5-1=0.5 so upstream speed is 0.5

OpenStudy (kitkat16):

downstream = 1.5+1=2.5

OpenStudy (kitkat16):

Need to find the speed in still water?

OpenStudy (kitkat16):

@jim_thompson5910 are u busy?

Directrix (directrix):

Upstream: 3 = (r - 1) * (t) where r is the boat speed in still water and t is the time spent rowing upstream. Downstream: 3 = (r + 1) * (4 - t) So, 3 = (r - 1) * (t) = (r + 1) * (4 - t) See what you think. It's just an idea.

Directrix (directrix):

So, (r - 1) * (t) = (r + 1) * (4 - t)

Directrix (directrix):

Which gives a rate of 2. I wish we could get somebody over here to check. I think that the notification features are malfunctioning on OS tonight.

OpenStudy (kitkat16):

yeah OS isnt working to well

Directrix (directrix):

Do you think 2 is incorrect? @KitKat16

OpenStudy (kropot72):

Still water speed of boat checks out at 2 miles per hour.

Directrix (directrix):

Thanks, @kropot72

OpenStudy (kitkat16):

Thanks guys for the help! :)

OpenStudy (kropot72):

You're welcome :)